# Linear Discriminant Regularized Regression

### v2

@@ 2.3 Classification based on reduced dimensions
we do not suffer numerically in solving (2.17), in stark contrast with the algorithmic issue of solving (1.6) (or a penalized version thereof) in the original $p$-dimensional space. As pointed out by Wu et al. (2015), the latter leads to meaningless solutions with the objective function equal to infinity. Finally, we remark that the requirement in (2.18) is easy to satisfy as ${\rm rank}({\mathbf{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbf{I}_{n}-P_{{\mathbf{Y}}})=n-L$ when we observe at least one data point per class.
3 Theoretical guarantees
In this section we provide a unified theory for analyzing the classifier $\widehat{g}$ in (2.13) based on estimating the regression coefficient matrix. In Section 3.1 we start with full generality and bound the excess risk of $\widehat{g}$ by a quantity that is related with the error of estimating $B^{*}$. The results there hold for any estimator $\widehat{B}^{*}$ of $B^{*}$. In Section 3.2 we focus on the proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ and provide a unified analysis of its estimation error that is valid for any generic estimator $\widehat{B}$. Finally we apply the general result to two particular choices of $\widehat{B}$, the lasso estimator in Section 3.3 and the reduced-rank estimator in Section 3.4.
3.1 A reduction scheme for bounding the excess risk
Let $(X,Y)$ be a new pair from (1.1) that is independent of the training data ${\bm{D}}:=\{{\mathbf{X}},{\mathbf{Y}}\}$. For any classifier $\widehat{g}:\mathbb{R}^{p}\to[L]$ in (2.13), its excess risk relative to the error rate of the Bayes rule in (1.2) is defined as
Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition. Our analysis of $\mathcal{R}(\widehat{g})$ requires the following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$.
Assumption 1.
There exists some absolute constants $0<c\leq C<\infty$ such that
This is a common regularity assumption even in binary classification problems. We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 21 of Appendix D) at the cost of more technical proofs and a less transparent presentation.
The following proposition states that for the purpose of bounding $\mathcal{R}(\widehat{g})$, it suffices to study the estimation error of each discriminant function, that is, $\widehat{G}_{\ell}(X)-G_{\ell}(X)$ for $\ell\in[L]$. The proof essentially follows the argument of proving Theorem 12 in Bing and Wegkamp (2023) but with modifications to capture the sum of squared errors of estimating all discriminant functions.
Proposition 2.
Under model (1.1) with 1, for any sequences $t_{1},\ldots,t_{L}$ (that possibly depend on ${\bm{D}}$), with probability equal to one, one has
$\mathcal{R}(\widehat{g})~{}\leq~{}C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\}.$
Proof.
See Section B.2. ∎
$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)$ (3.2)
for each $\ell\in[L]$. It is worth mentioning that in this section we allow $\widehat{B}^{*}$ in the above display to be any estimator of $B^{*}$ whereas $\widehat{\mu}_{\ell}$ and $\widehat{\pi}_{\ell}$ are given in (2.9). Since the variance of the first term in (3.2) depends on $\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$, we define
$\Delta_{\infty}:=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under $\mathbb{E}(X)=0_{p}$)
${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}).$
1, the Gaussian tail of $X\mid Y$ and the independence between $X$ and $({\mathbf{X}},{\mathbf{Y}})$ in (3.2) entail the following upper bound of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
Theorem 3.
Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$ and $L\log(n)\leq n$. With probability at least $1-3n^{-1}$, we have, for all $\ell\in[L]$,
$\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|~{}\lesssim~{}\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}$ $+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}$
$+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
Proof.
See Section B.3. ∎
Condition $L\log(n)\leq n$ together with 1 ensures that $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$. The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish. It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
$\mathcal{Q}(\widehat{B}^{*}):=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}(\Delta_{\infty}+\log n)+\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}.$ (3.3)
Corollary 4.
Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$. For any estimator $\widehat{B}^{*}$ of $B^{*}$, the corresponding classifier $\widehat{g}$ in (2.13) satisfies
$\mathbb{E}_{\bm{D}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\mathbb{E}_{{\bm{D}}}\left[\min\left\{{\Delta_{\infty}}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right].$
Proof.
See Section B.4. ∎
4 ensures that bounding $\mathcal{R}(\widehat{g})$ can be reduced to controlling $\mathcal{Q}(\widehat{B}^{*})$ for any estimator $\widehat{B}^{*}$ of $B^{*}$. In particular, 4 is valid for any existing approach based on estimating $B^{*}$. In the next section we provide a unified analysis of $\mathcal{Q}(\widehat{B}^{*})$ for our proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ based on a generic $\widehat{B}$. As shown later, the first term in (3.3) is typically the dominating term.
Assumption 2.
We need the following regularity conditions on $\Sigma_{w}$. It assumes the largest entry of $\Sigma_{w}$ is bounded from above and the smallest eigenvalue of $\Sigma_{w}$ is bounded away from zero.
Assumption 3.
There exists some positive constants $0<c\leq C<\infty$ such that $c<\lambda_{p}(\Sigma_{w})\leq\|\Sigma_{w}\|_{\infty}\leq C$.
Theorem 5.
Under model (1.1) with Assumptions 1 & 3, assume $\Delta\asymp 1$ and $L\log(n)\leq cn$ for some small constant $c>0$. For any $\widehat{B}$ satisfying 2 with $\omega_{2}\sqrt{L}\leq c,$ with probability at least $1-2n^{-1}$,
the matrix $\widehat{H}$ in (2.10) is invertible,
and
Proof.
See Section B.5. ∎
Theorem 5 relates $\mathcal{Q}(\widehat{B}^{*})$ to the estimation error of $\widehat{B}$ in both the Frobenius norm and the $\ell_{1}/\ell_{2}$-norm. In the bound of (3.4), the second term on the right-hand-side is usually the leading term and its rate depends on specific choice of $\widehat{B}$. We provide concrete examples in Section 3.3 for the lasso estimator and in Section 3.4 for the reduced-rank estimator.
Condition $\omega_{2}\sqrt{L}\leq c$ ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds so that $\widehat{H}$ is invertible. Furthermore, the bound in (3.4) suggests that the condition $\omega_{2}\sqrt{L}\leq c$ is also needed for consistent classification.
3.3 Application to estimation of $B$ with entry-wise $\ell_{1}$-regularization
For ease of presentation, we simply use the largest number of non-zero entries among all columns of $B$. However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$. Our analysis also needs the following mild regularity condition on the matrix ${M}$ of conditional means.
Assumption 4.
The following theorem establishes explicit rates of $\omega_{2}$ and $\omega_{1}$ in 2 for the regression matrix estimator $\widehat{B}$ in (1.14) using the lasso-penalty in (2.7).
Theorem 6.
Under model (1.1) with Assumptions 1, 3 and 4, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with $(s\vee L)\log(n\vee p)\leq c~{}n.$ Let the estimator $\widehat{B}$ be obtained from (1.14) with any $\lambda$ satisfying
and some finite, large enough constant $C>0$. Then, with probability at least $1-3n^{-1}$, 2 holds for
$\omega_{2}\asymp\lambda\sqrt{sL},\qquad\omega_{1}\asymp\lambda sL.$
Proof.
See Section B.6. ∎
Our proof is in fact based on a weaker condition than 3 by replacing $\lambda_{p}(\Sigma_{w})$ with the Restricted Eigenvalue condition (RE) on $\Sigma_{w}$. The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$. For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbf{Y}}$ is not a linear model in ${\mathbf{X}}$. We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$ in Lemma 11. For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbf{X}}$ are generated from a mixture of Gaussians. In Lemma 12 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b). Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}^{\prime}-{\mathbf{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbf{Y}}^{\prime}$ and different target matrix $B^{\prime}$. Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
In the rest of the paper, we assume that $\lambda$ is chosen as the order specified in (3.5). In practice, one can simply choose $\lambda$ via cross-validation, for instance, via the built-in function of the R-package glmnet.
Corollary 7.
Under the conditions in Theorem 6, assume that $\Delta_{\infty}\asymp 1$. Then
Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$. The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$. Guarantees on the excess risk are stronger since they together with the Bayes error, which is oftentimes easy to calculate, imply guarantees of the risk.
7 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$. Otherwise, if the pairwise separation between different classes grows to infinity, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)). Indeed, provided $\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$ using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}{sL^{2}\over n}\log^{2}(n\vee p)e^{-c\Delta_{\infty}}+n^{-C_{0}}$
for some arbitrary large constant $C_{0}>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
3.4 Application to estimation of $B$ via reduced-rank regression
We provide analysis for the classifier that estimates the regression coefficient matrix $B$ by reduced-rank regression. Specifically, let $\widehat{B}$ be obtained from (1.14) with $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ for some tuning parameter $\lambda>0$. The following theorem establishes the rates of $\omega_{2}$ and $\omega_{1}$ for which 2 holds. Write $r={\rm rank}(B)$.
Theorem 8.
Under model (1.1) with Assumptions 1 and 3, assume $(p+L)\log(n)\leq c~{}n$ for some sufficiently small constant $c>0$. For the estimator $\widehat{B}$ in (1.14) with any
$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}$ (3.6)
and some finite, large enough constant $C>0$, with probability at least $1-2n^{-1}$, 2 holds for
Proof.
The proof is deferred to Section B.7. ∎
For the reduced-rank estimator, establishing its in-sample prediction risk can be done without any condition on the design matrix (see, for instance, Bunea et al. (2011); Giraud (2011)). However, since 2 is also related with the out-of-sample prediction risk, we need the smallest eigenvalue of the Gram matrix $\widehat{\Sigma}$ to be bounded away from zero. The inequality $(p+L)\log(n)\leq cn$ is assumed for this purpose. In case $p$ is large, it is recommended that an additional ridge penalty be added in $\text{pen}(B)$ to alleviate the singularity issue of $\widehat{\Sigma}$ (Mukherjee and Zhu, 2011). Another main ingredient of our proof of Theorem 8 is to control $\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}$ in Lemma 13. It requires non-standard analysis for the same reason mentioned before that ${\mathbf{Y}}$ is not linearly related with ${\mathbf{X}}$.
Corollary 9.
Under the conditions in Theorem 8, assume $\Delta_{\infty}\asymp 1$ and $\|\Sigma_{w}\|_{\rm op}\lesssim 1$. Then
Note that for the $p\times L$ matrix $B$ with rank $r$, its effective number of parameters is exactly $(p+L)r$ (Izenman, 1975) which could be much smaller than $pL$ when $r\ll(p\wedge L)$. 9 suggests that the classifier using the reduced-rank estimator has benefit when $r$ is relatively small comparing to $p$ and $L$.
The same remark as the one at the end of Section 3.3 applies, and our analysis can be extended to $\Delta_{\infty}\to\infty$, where the rate in 9 becomes exponentially fast in $\Delta_{\infty}$.
4 Simulation study
In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR). We also include the procedure in Section 2.3 that is based on Fisher’s discriminant rule after estimating $B$ (called LDRR-F). For estimating $B$ from (1.14), in the sparse scenarios of Section 4.1, we consider both the lasso penalty (2.7) (L1) and the elastic net penalty (L1+L2), $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{1}+(1-\alpha)/2\|B_{\ell}\|_{2}^{2}]$, for some $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet. In the low-rank scenarios of Section 4.2, we consider both the reduced-rank penalty (2.8) (RR) and the reduced-rank penalty plus the ridge penalty (RR+L2), $\text{pen}(B)=\lambda[\alpha~{}{\rm rank}(B)+(1-\alpha)/2\|B\|_{F}^{2}]$, with tuning parameters $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation.
We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al. (2002), the shrunken centroids regularized discriminant analysis (RDA) method of Guo et al. (2007) and the $\ell_{1}$-penalized linear discriminant (PenalizedLDA) method of Witten and Tibshirani (2011). These methods are available in R-packages pamr, rda, penalizedLDA, and we use their default methods for selecting tuning parameters. Finally, the oracle procedure that uses the true values of ${M}$, $\Sigma_{w}$, and $\pi$ in computing the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.
@@ 4.1 Sparse scenarios
We generate the data from the following mean-shift scenario. A similar setting is also used in Witten and Tibshirani (2011). Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $[\mu_{\ell}]_{j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0. Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$. The quantity $\sigma>0$ controls the overall noise level while the coefficient $\rho\in[0,1]$ controls the within-class correlation among the features. We generate the class probabilities $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ as follows. For a given $\alpha\geq 0$, we set $\pi_{\ell}=\nu_{\ell}^{\alpha}/(\sum_{i=1}^{L}\nu_{i}^{\alpha})$ for all $\ell\in[L]$. Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform$(0,1)$. The quantity $\alpha$ controls the imbalance of each class size: a larger $\alpha$ corresponds to more imbalanced class sizes. At the other extreme, $\alpha=0$ corresponds to the balanced case $\pi=(1/L)1_{L}$. In the following, we examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{100,300,500,700,900\}$; the feature dimension $p\in\{100,300,500,700,900\}$; the number of classes $L\in\{2,5,8,11,14\}$; the correlation coefficient $\rho\in\{0,0.2,0.4,0.6,0.8\}$; the noise parameter $\sigma\in\{0.6,0.9,1.2,1.5,1.8\}$; the imbalance of the class probabilities $\alpha\in\{0,0.4,0.8,1.2,1.6\}$.
@@
From Figure 1, it is clear that our procedures consistently outperform other methods across all settings. We observe that the performance of our algorithms improves as $n$ increases, $L$ decreases and $\sigma$ decreases. Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms. These findings are in line with our theory in Section 3.3. It is a bit surprising to see that the classification error of our benchmark decreases as the correlation coefficient $\rho$ keeps increasing after $0.4$. Nevertheless, both our algorithms and RDA adapt to this situation. Finally, we find that both LDRR and LDRR-F as well as their lasso and elastic net variants have very comparable performance in all scenarios. This is as expected as the data is simulated according to model (1.1), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
Figure 1: The averaged misclassification errors of each algorithm for varying each parameter one at a time.

### v3

@@ 4.1 A practical classification based on reduced dimensions
we do not encounter numerical difficulties when solving (4.4), in stark contrast with the algorithmic issue of solving (1.4) (or a penalized version thereof) in the original $p$-dimensional space. As pointed out by Wu et al. (2015), the latter leads to meaningless solutions with the objective function equal to infinity. Finally, we remark that the requirement in (4.5) is easy to satisfy as ${\rm rank}({\mathbf{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbf{I}_{n}-P_{{\mathbf{Y}}})=n-L$ when we observe at least one data point per class.
@@ 4.2 Sparse scenarios
We generate the data from the following mean-shift scenario. A similar setting is also used in Witten and Tibshirani (2011). Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $\mu_{\ell}]_{j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0. Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$. The quantity $\sigma>0$ controls the overall noise level while the coefficient $\rho\in[0,1]$ controls the within-class correlation among the features. We generate the class probabilities $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ as follows. For a given $\alpha\geq 0$, we set $\pi_{\ell}=\nu_{\ell}^{\alpha}/(\sum_{i=1}^{L}\nu_{i}^{\alpha})$ for all $\ell\in[L]$. Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform(0,1)(0,1). The quantity $\alpha$ controls the imbalance of each class size: a larger $\alpha$ corresponds to more imbalanced class sizes. At the other extreme, $\alpha=0$ corresponds to the balanced case $\pi=(1/L)1_{L}$. In the following, we examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{100,300,500,700,900\}$; the feature dimension $p\in\{100,300,500,700,900\}$; the number of classes $L\in\{2,5,8,11,14\}$; the correlation coefficient $\rho\in\{0,0.2,0.4,0.6,0.8\}$; the noise parameter $\sigma\in\{0.6,0.9,1.2,1.5,1.8\}$; the imbalance of the class probabilities $\alpha\in\{0,0.4,0.8,1.2,1.6\}$.
@@
From Figure 1, it is clear that our procedures consistently outperform other methods across all settings. The performance of our algorithms improves as $n$ increases, $L$ decreases and $\sigma$ decreases. Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms. These findings are in line with our theory in Section 3.3. It is a bit surprising to see that the classification error of our benchmark decreases as the correlation coefficient $\rho$ keeps increasing after $0.4$. Nevertheless, both our algorithms and RDA adapt to this situation. Finally, we find that both LDRR and LDRR-F as well as their lasso and elastic net variants have very comparable performance in all scenarios. This is as expected as the data is simulated according to model (3.1), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
Figure 1: The averaged misclassification errors in sparse scenarios.